# Tree Search for Language Model Agents

### v2

@@ 4.2 Results
SteP (Sodhi et al., 2024) 35.8% - -
@@
Our results are summarized in Tab. 2. Introducing search increases success rate substantially across the board. Search improves the success rate of the baseline GPT-4o + SoM agent on VWA by 39.7% relatively (increasing from 18.6% to 26.4%), setting a new state-of-the-art on the benchmark. On WA, introducing search to the GPT-4o agent improves the success rate substantially as well, increasing it by 28.0% relatively (15.0% to 19.2%). This is competitive with other prompt-based agents on WA, but in future work it will be interesting to explore introducing search to stronger baseline agents that incorporate domain-specific techniques, such as SteP (Sodhi et al., 2024) or AutoGuide (Fu et al., 2024).
@@ 5.1 Ablations
We conduct several ablation experiments on a subset of 200 tasks from VWA (100 Shopping tasks, 50 Reddit task, and 50 Classifieds tasks).
@@
We plot the success rate of the GPT-4o agent with search limited to varying budgets $c\in\{0,5,10,15,20\}$ in Fig. 2. All experiments are conducted with search parameters of depth $d=5$ and branching factor $b=5$. The search budget specifies the maximum number of node expansions performed at each step. For example, a search budget of 10 indicates that at most 10 nodes will be expanded, after which the agent will commit to and execute the trajectory with the highest value. We observe that success rate generally increases as search budget increases. Notably, performing even very small amounts of search ($c=5$) substantially improves success rate by 30.6% relative to not doing search (24.5% to 32.0%). When the budget is increased to $c=20$, this improves success rate by 51.0% relative to not doing search (from 24.5% to 37.0%), highlighting the benefit of scaling the search budget. Running experiments with an even greater search budget to evaluate scaling trends would be a promising future direction to explore.
@@
We run an ablation experiment varying the search branching factor $b$ and maximum depth $d$. The results are summarized in Tab. 3. We observe that in general, success rate increases as the size of search tree increases (along both $b$ and $d$ dimensions). In particular, scaling both $b$ and $d$ is necessary to achieve strong performance.
@@
We ablate the multimodal model used for the value function, swapping out GPT-4o for (1) the LLaVA-v1.6-34B (Liu et al., 2024a) multimodal model prompted zero-shot (with just the current observation) and (2) the groundtruth reward from VWA (which is a sparse reward signal that returns either 0 or 1), and (3) GPT-4o without self-consistency. The results are summarized in Tab. 4. We find that the GPT-4o value function significant outperforms the LLaVA model, improving the result of the agent from 30.0% to 37.0%. The groundtruth reward function achieves a success rate of 43.5%. These results suggest that there is still significant headroom in improving the search algorithm by training better value functions. We also observe that self-consistency is essential for the performance of search ($28.5\%\rightarrow 37.0\%$), which we attribute to it enabling marginalization over multiple reasoning chains, reducing noise during state evaluation.
@@ 5.2 Success Rate Breakdown
The VWA benchmark includes labels for the action difficulty of each task. These labels are human annotated, and roughly indicate the number of actions a human would need to take to solve the tasks: easy tasks require 3 or fewer actions, medium tasks require 4–9 actions, and hard tasks demand 10 or more. These guidelines are approximate and devised by the human annotators of VWA, so there may exist more optimal solutions in practice. The increase in success rate from introducing search is summarized in Tab. 5. Introducing search improves performance across all difficulty levels, but it introduces much larger gains on tasks of medium action difficulty, with a relative increase of 75% in success rate (from 12.7% to 22.2%). We hypothesize that this is because our search parameters (max depth $d=5$) are beneficial for a large proportion of medium difficulty tasks. Conversely, achieving even better performance on hard tasks may require search over deeper trees. Easy tasks likely do not benefit as much from search, as they generally involve less multi-step planning (some can be solved with 1 or 2 actions), and baselines already have higher success rates.
Figure 3: Search can improve the robustness of agents by filtering out bad actions. Shown above is a trajectory for VWA classifieds task #48 where greedily picking the first sampled actions would have led to a failure (by taking the path in the first row). Search avoids this failure mode by exploring and pruning less promising paths, ultimately committing to the highlighted trajectory.
@@
Tables 7 and 7 summarize the success rates across the various websites in the VWA and WA benchmarks. We observe an improvement in success rates across the board, demonstrating that our method generalizes across sites. Specifically, the increase is most substantial on the Classifieds and Shopping sites in VWA, with relative increases of 44% and 45%, respectively. Similarly, the CMS site in the WA benchmark shows a significant relative improvement of 50%.
@@ 5.3 Qualitative Results
Many tasks in VWA and WA require an agent to keep a persistent memory of multiple previous actions and observations. A common failure mode amongst agents without search is that they tend to undo previous actions, or get stuck in loops (see Appendix C.4 of Koh et al. 2024). An example for VWA shopping task #256 is shown in Fig. 1, where the agent is tasked to add two different types of canned fruit from the same brand to the comparison list. The baseline agent successfully adds the first item, but fails to navigate to the second item, as it returns to the homepage in step 3 and gets confused. This is an example of compounding error leading to overall task failure, which is fairly common in existing baseline agents without search.
When search is introduced, the agent explores other plausible trajectories and backtrack when those eventually result in failure: the same GPT-4o agent with search is able to find a successful multi-step trajectory for the same task, which involves adding the first item (action #1 in Fig. 1), typing in a search query (action #6), and adding the correct second item to the comparison list (action #9).
@@
An inherent issue with sampling actions from language models is that we are sampling from a distribution over text, and the first sample we generate may not always be the best action to take in the environment. Search allows us to evaluate each generated action concretely by executing it in the simulator, and use the received environmental feedback to make better decisions. One example is VWA classifieds task #48 (Fig. 3), which is to find a post containing a particular image. If the agent executes the first sampled action at every step (i.e., the sequence in the top row), it results in failure. Search allows the agent to enumerate all possibilities by executing plausible actions and receiving environment feedback, which allows it to select the best (and in this case, successful) trajectory.
@@ 5.4 Limitations
While we have shown that introducing search to language model agents achieves promising results on web tasks, it does come with some practical limitations:
@@
Introducing search allows us to expend more compute at inference time to extract stronger results from the baseline LM agent. However, this results in trajectories taking significantly longer to execute, as the agent has to perform more exploration and hence more inference calls to the LM. For example, a search budget of $c=20$ implies that an agent with search could potentially expand up to 20 states in each search iteration, which would use up to 20×20$\times$20 × more LM calls than an agent without search. Research on improving the efficiency and throughput of machine learning systems (Leviathan et al., 2023; Dao et al., 2022; Dao, 2023) will likely help with optimizing this, but for practical deployment one may need to carefully set the search parameters $b$, $d$, and $c$ to balance between achieving better results and overall time spent completing a task.
In our approach, we implemented search by keeping track of the sequence of actions required to get to a state. During backtracking, we reset the environment and apply the same sequence after resetting the environment. This is necessary, as naively executing the go_back action (Tab. 1) may discard important information on the page, such as the scroll offset and already entered text. However, these environment calls for backtracking introduce additional overhead, which may be restrictive for deployment if calls to the environment are expensive.
@@
For real world deployment, we will need to restrict the search space to actions that are not destructive. Destructive actions are defined as actions that will irreversibly change the state of the website and are difficult to backtrack from. For example, placing an order on an e-commerce site is typically very difficult to automatically undo. One way to address this is to introduce a classifier that predicts when certain actions are destructive, and prevent node expansion for those states. If we have specific domain knowledge about the downstream application (e.g., we know certain pages should be off limits), such rules can be manually enforced with high accuracy. One advantage of tree search is that it is easier to incorporate such a constraint: it can be directly integrated into the value function to prevent execution of dangerous actions. Another direction to handle this would be to train a world model (Ha & Schmidhuber, 2018) that we can use for simulations during search. Search may also be more easily implemented in offline settings where actions are non-destructive as they can always be undone or reset, such as programming (Jimenez et al., 2023; Yang et al., 2024) or Microsoft Excel (Li et al., 2024).
@@ 6 Conclusion
In this paper, we introduced an inference-time search algorithm designed to enhance the capabilities of language model agents on realistic web tasks. Our approach integrates best-first tree search with LM agents, enabling them to explore and evaluate multiple action trajectories to achieve superior performance on web tasks. This is the first time search has been shown to significantly improve the success rates of LM agents on realistic web environments, as demonstrated on the (Visual)WebArena benchmarks. Our search procedure is general, and it will be valuable to apply it to other domains in future work. We believe that inference-time search will be a key component for building capable agents that can plan, reason, and act autonomously to perform computer tasks.
Acknowledgments
We thank Andy Liu, Saujas Vaduguru, Zhiruo Wang, Alex Xie, Shuyan Zhou, and many others for helpful feedback and discussions on previous versions of this paper, and Santiago Cortes-Gomez for infrastructure support. We thank Wendy Kua for helping with the figures.
Ethics Statement
As an active area of machine learning research, language model web agents present both opportunities and potential ethical considerations. Improved web agents could improve accessibility for users with disabilities, automate repetitive or tedious tasks, and potentially democratize access to complex web platforms. Our search method contributes towards making such benefits more reliable and widely available by improving the robustness and success rate of language model agents. However, we acknowledge several potential ethical considerations:
@@ Acknowledgments
Reproducibility Statement
To ensure reproducibility of our results, we provide details of the tree search algorithm within this paper (Sec. 3.3) as well as more formally in Appendix. A.4. The details of the models used are also provided in Sec. 4.1. All other prompts and implementation details necessary to reproduce our results are provided in Appendix A.3. In order to ensure long term reproducibility, we also provide results for a setting with open model weights for longer term reproducibility, the Llama-3 agent with the LLaVA value function (Appendix. A.2.1), in addition to the API-based agents (which at present achieve higher performance than open sourced alternatives). All of our code and experiment launch scripts are open sourced and made publicly available on GitHub at jykoh.com/search-agents.
@@ A.2.1 Value Function Ablations
A.2.2 Comparison to Trajectory-Level Reranking
We observe that this re-ranking baseline starts to plateau around 7 runs, which achieves a success rate of 30%. This underperforms our approach with search budget $c\geq 5$ (Fig. 2). It is also substantially worse than our approach with $c=20$, which achieves a success rate of 37.0% on the ablation subset.
@@ A.3.1 Language Model Agents
You are an autonomous intelligent agent tasked with navigating a web browser. You will be given web-based tasks. These tasks will be accomplished through the use of specific actions you can issue. Here’s the information you’ll have: The user’s objective: This is the task you’re trying to complete. The current web page screenshot: This is a screenshot of the webpage, with each interactable element assigned a unique numerical id. Each bounding box and its respective id shares the same color. The observation, which lists the IDs of all interactable elements on the current web page with their text content if any, in the format [id] [tagType] [text content]. tagType is the type of the element, such as button, link, or textbox. text content is the text content of the element. For example, [1234] [button] [’Add to Cart’] means that there is a button with id 1234 and text content ’Add to Cart’ on the current web page. [] [StaticText] [text] means that the element is of some text that is not interactable. The current web page’s URL: This is the page you’re currently navigating. The open tabs: These are the tabs you have open. The previous action: This is the action you just performed. It may be helpful to track your progress. The actions you can perform fall into several categories: Page Operation Actions: ```click [id]```: This action clicks on an element with a specific id on the webpage. ```type [id] [content]```: Use this to type the content into the field with id. By default, the “Enter” key is pressed after typing unless press_enter_after is set to 0, i.e., ```type [id] [content] [0]```. ```hover [id]```: Hover over an element with id. ```press [key_comb]```: Simulates the pressing of a key combination on the keyboard (e.g., Ctrl+v). ```scroll [down]``` or ```scroll [up]```: Scroll the page up or down. Tab Management Actions: ```new_tab```: Open a new, empty browser tab. ```tab_focus [tab_index]```: Switch the browser’s focus to a specific tab using its index. ```close_tab```: Close the currently active tab. URL Navigation Actions: ```goto [url]```: Navigate to a specific URL. ```go_back```: Navigate to the previously viewed page. ```go_forward```: Navigate to the next page (if a previous ’go_back’ action was performed). Completion Action: ```stop [answer]```: Issue this action when you believe the task is complete. If the objective is to find a text-based answer, provide the answer in the bracket. Homepage: If you want to visit other websites, check out the homepage at http://homepage.com. It has a list of websites you can visit. http://homepage.com/password.html lists all the account name and password for the websites. You can use them to log in to the websites. To be successful, it is very important to follow the following rules: 1. You should only issue an action that is valid given the current observation 2. You should only issue one action at a time. 3. You should follow the examples to reason step by step and then issue the next action. 4. Generate the action in the correct format. Start with a “In summary, the next action I will perform is” phrase, followed by action inside ``````. For example, “In summary, the next action I will perform is ```click [1234]```”. 5. Issue stop action when you think you have achieved the objective. Don’t generate anything after stop.

### v3

@@ 4.2 Results
SteP (Sodhi et al., 2024) 33.5% - -
GUI-API Hybrid Agent (Song et al., 2024) 35.8% - -
AgentOccam (Yang et al., 2024b) 43.1% - -
AgentOccam-Judge (Yang et al., 2024b) 45.7% - -
@@
Our results are summarized in Tab. 2. Introducing search increases success rate substantially across the board. Search improves the success rate of the baseline GPT-4o + SoM agent on VWA by 39.7% relatively (increasing from 18.9% to 26.4%), setting a new state-of-the-art on the benchmark. On WA, introducing search to the GPT-4o agent improves the success rate substantially as well, increasing it by 28.0% relatively (15.0% to 19.2%). While other baseline agents obtain higher performance than our GPT-4o baseline through domain-specific techniques such as introducing website specific guidelines (Sodhi et al., 2024; Fu et al., 2024) or engineering improved input spaces (Song et al., 2024; Yang et al., 2024b), these techniques are orthogonal to — and potentially complementary with — our search-based approach.
@@ 5.1 Ablations
We conduct several ablations on a subset of 200 VWA tasks.
@@
We plot the success rate of the GPT-4o agent with search limited to varying budgets $c\in\{0,5,10,15,20\}$ in Fig. 2. All experiments are conducted with search parameters of depth $d=5$ and branching factor $b=5$. The search budget specifies the maximum number of node expansions performed at each step. For example, a search budget of 10 indicates that at most 10 nodes will be expanded, after which the agent will commit to and execute the trajectory with the highest value. We observe that success rate generally increases as search budget increases. Notably, performing even very small amounts of search ($c=5$) substantially improves success rate by 30.6% relative to not doing search (24.5% to 32.0%). When the budget is increased to $c=20$, this improves success rate by 51.0% relative to not doing search (from 24.5% to 37.0%), highlighting the benefit of scaling the search budget.
@@
We run an ablation experiment varying the search branching factor $b$ and maximum depth $d$. The results are summarized in Tab. 3. We observe that in general, success rate increases as the size of search tree increases (along both $b$ and $d$ dimensions), and scaling both $b$ and $d$ is necessary to achieve strong performance.
@@
We ablate the multimodal model used for the value function, swapping out GPT-4o for (1) the LLaVA-v1.6-34B (Liu et al., 2024a) multimodal model prompted zero-shot (with only the current observation, as LLaVA only supports a single image input) and (2) the groundtruth reward from VWA (which is a sparse reward signal that returns either 0 or 1, and does not track partial progress), and (3) GPT-4o without self-consistency. The results are summarized in Tab. 4. We find that the GPT-4o value function significant outperforms the LLaVA model, improving the result of the agent from 30.0% to 37.0%. The groundtruth reward function achieves a success rate of 43.5%. These results suggest that there is still significant headroom in improving the search algorithm with better value functions. We also observe that self-consistency is essential for good performance ($28.5\%\rightarrow 37.0\%$), which we attribute to it enabling marginalization over multiple reasoning chains, reducing noise during state evaluation. While we do not finetune custom value function models in this paper, we believe that this is a promising direction for future work, and may result in value function models that can outperform our current version.
Comparison to Trajectory-Level Reranking
We observe that this re-ranking baseline starts to plateau around 7 runs, which achieves a success rate of 30%. This underperforms our approach with search budget $c\geq 5$ (Fig. 2). For $c=5$ and $n=5$, the agent from these two approaches spend approximately equal inference compute. It is also substantially worse than our approach with $c=20$, which achieves a success rate of 37.0% on the ablation subset.
@@ 5.2 Success Rate Breakdown
The VWA benchmark includes labels for the action difficulty of each task. These labels are human annotated, and roughly indicate the number of actions a human would need to take to solve the tasks: easy tasks require 3 or fewer actions, medium tasks require 4–9 actions, and hard tasks demand 10 or more. These guidelines are approximate and devised by the human annotators of VWA, so there may exist more optimal solutions in practice. The increase in success rate from introducing search is summarized in Tab. 5. Introducing search improves performance across all difficulty levels, but introduces much greater gains in medium difficulty tasks, with a relative increase of 75% in success rate (from 12.7% to 22.2%). We hypothesize that this is because our search parameters (max depth $d=5$) are beneficial for a large proportion of medium difficulty tasks. Conversely, achieving even better performance on hard tasks may require search over deeper trees. Easy tasks do not benefit as much from search, as they generally involve less planning (some can be solved with 1 or 2 actions), and baselines already have higher success rates.
Figure 4: Search can improve robustness by backtracking from bad actions. Shown above is a trajectory for VWA classifieds task #48 where greedily picking the first sampled actions would have led to a failure (the path in the first row).
@@
Tables 7 and 7 summarize the success rates across the various websites in the VWA and WA benchmarks. We observe an improvement in success rates across the board, demonstrating that our method generalizes across sites. Specifically, the increase is most substantial on the Classifieds and Shopping sites in VWA, with relative increases of 44% and 45%, and the CMS site in the WA benchmark (relative improvement of 50%).
@@ 5.3 Qualitative Results
Many tasks in VWA and WA require an agent to keep a persistent memory of multiple previous actions and observations. A common failure mode amongst agents without search is that they tend to undo previous actions, or get stuck in loops (see Appendix C.4 of Koh et al. 2024). An example for VWA shopping task #256 is shown in Fig. 1, where the agent is tasked to add two different types of canned fruit from the same brand to the comparison list. The baseline agent successfully adds the first item, but fails to navigate to the second item, as it returns to the homepage in step 3 and gets confused. This is an example of compounding error leading to overall task failure, which is fairly common in existing baseline agents without search. When search is introduced, the agent explores other plausible trajectories and backtracks when those eventually result in failure: the same GPT-4o agent with search is able to find a successful multi-step trajectory for the same task, which involves adding the first item (action #1 in Fig. 1), typing in a search query (#6), and adding the correct second item to the comparison list (#9).
@@
An inherent issue with sampling actions from language models is that we are sampling from a distribution over text, and the first sample we generate may not always be the best action to take in the environment. Search allows us to evaluate each generated action concretely by executing it in the simulator, and use the received environmental feedback to make better decisions. One example is VWA classifieds task #48 (Fig. 4), which is to find a post containing a particular image. If the agent executes the first sampled action at every step (i.e., the sequence in the top row), it results in failure. Search allows the agent to explore and enumerate multiple possibilities.
@@ 5.4 Limitations
While we have shown that introducing search to language model agents achieves promising results on web tasks, our approach does come with some practical considerations.
@@
Introducing search allows us to expend more compute at inference time to extract stronger results from the baseline LM agent. However, this results in trajectories taking significantly longer to execute, as the agent has to perform more exploration and hence more inference calls to the LM. For example, a search budget of $c=20$ implies that an agent with search could potentially expand up to 20 states in each search iteration, which would use up to $20\times$ more LM calls than an agent without search. Research on improving the efficiency and throughput of machine learning systems (Leviathan et al., 2023; Dao et al., 2022; Dao, 2023) will likely help with optimizing this, but for practical deployment one may need to carefully set the search parameters $b$, $d$, and $c$ to balance between achieving better results and overall time spent completing a task.
In our approach, we implemented search by keeping track of the sequence of actions required to get to a state. During backtracking, we reset the environment and apply the same sequence after resetting the environment. This is necessary, as naively executing the go_back action (Tab. 1) may discard important information on the page, such as the scroll offset and already entered text.
@@
For real world deployment, we will need to restrict the search space to actions that are not destructive. Destructive actions are defined as actions that will irreversibly change the state of the website and are difficult to backtrack from. For example, placing an order on an e-commerce site is typically difficult to undo. One way to address this is to introduce a classifier that predicts when certain actions are destructive, and prevent node expansion for those states. If we have specific domain knowledge about the downstream application (e.g., we know certain pages should be off limits), such rules can be manually enforced with high accuracy. One advantage of tree search is that it is easier to incorporate such a constraint: it can be directly integrated into the value function to prevent execution of dangerous actions. Another direction to handle this would be to train a world model (Ha & Schmidhuber, 2018) that we can use for simulations during search. Search may also be more easily implemented in offline settings where actions are non-destructive as they can always be undone or reset, such as programming (Jimenez et al., 2023; Yang et al., 2024a) or Microsoft Excel (Li et al., 2024).
Domain Specific Value Functions
In this work, we only consider tree search in the context of agents operating in dynamic web environments, and our value function is tailored to capture the nuances of web navigation, incorporating features such as page content and navigation history (in the form of previously seen screenshots). This is likely to be less effective for other agentic domains, such as software engineering (SWE) agents Yang et al. (2024a). Although our overall tree search approach is generalizable, the value function may require adaptation to effectively incorporate domain-specific context: for example, code execution traces for SWE tasks. We leave detailed explorations for future work.
@@ 6 Conclusion
In this paper, we introduced an inference-time search algorithm designed to enhance the capabilities of language model agents on realistic web tasks. Our approach integrates best-first tree search with LM agents, enabling them to explore and evaluate multiple action trajectories to achieve superior performance on web tasks. This is the first time search has been shown to significantly improve the success rates of LM agents on realistic web environments, as demonstrated on the (Visual)WebArena benchmarks. Our search procedure is general, and it will be valuable to apply it to other domains in future work, or incorporate more sophisticated search algorithms such as MCTS. We believe that inference-time search will be a key component for building capable agents that can plan, reason, and act autonomously to perform computer tasks.
Statement of Broader Impact
As an active area of machine learning research, language model web agents present both opportunities and potential ethical considerations. Improved web agents could improve accessibility for users with disabilities, automate repetitive or tedious tasks, and potentially democratize access to complex web platforms. Our search method contributes towards making such benefits more reliable and widely available by improving the robustness and success rate of language model agents. However, we acknowledge several considerations of broader impact:
@@
  • Privacy: More capable web agents could potentially be used to scrape personal information or navigate private areas of websites more effectively. We emphasize the importance of respecting user privacy and website terms of service in any real-world deployment of these technologies.
@@ A.3.1 Language Model Agents
You are an autonomous intelligent agent tasked with navigating a web browser. You will be given web-based tasks. These tasks will be accomplished through the use of specific actions you can issue.
Here’s the information you’ll have:
The user’s objective: This is the task you’re trying to complete.
The current web page screenshot: This is a screenshot of the webpage, with each interactable element assigned a unique numerical id. Each bounding box and its respective id shares the same color.
The observation, which lists the IDs of all interactable elements on the current web page with their text content if any, in the format [id] [tagType] [text content]. tagType is the type of the element, such as button, link, or textbox. text content is the text content of the element. For example, [1234] [button] [’Add to Cart’] means that there is a button with id 1234 and text content ’Add to Cart’ on the current web page. [] [StaticText] [text] means that the element is of some text that is not interactable.
The current web page’s URL: This is the page you’re currently navigating.
The open tabs: These are the tabs you have open.
The previous action: This is the action you just performed. It may be helpful to track your progress.
The actions you can perform fall into several categories:
Page Operation Actions:
```click [id]```: This action clicks on an element with a specific id on the webpage.
```type [id] [content]```: Use this to type the content into the field with id. By default, the “Enter” key is pressed after typing unless press_enter_after is set to 0, i.e., ```type [id] [content] [0]```.
```hover [id]```: Hover over an element with id.
```press [key_comb]```: Simulates the pressing of a key combination on the keyboard (e.g., Ctrl+v).
```scroll [down]``` or ```scroll [up]```: Scroll the page up or down.
Tab Management Actions:
```new_tab```: Open a new, empty browser tab.
```tab_focus [tab_index]```: Switch the browser’s focus to a specific tab using its index.
```close_tab```: Close the currently active tab.
URL Navigation Actions:
```goto [url]```: Navigate to a specific URL.
```go_back```: Navigate to the previously viewed page.
```go_forward```: Navigate to the next page (if a previous ’go_back’ action was performed).
Completion Action:
```stop [answer]```: Issue this action when you believe the task is complete. If the objective is to find a text-based answer, provide the answer in the bracket.
Homepage:
If you want to visit other websites, check out the homepage at http://homepage.com. It has a list of websites you can visit.
http://homepage.com/password.html lists all the account name and password for the websites. You can use them to log in to the websites.
To be successful, it is very important to follow the following rules:
1. You should only issue an action that is valid given the current observation
2. You should only issue one action at a time.
3. You should follow the examples to reason step by step and then issue the next action.
4. Generate the action in the correct format. Start with a “In summary, the next action I will perform is” phrase, followed by action inside ``````. For example, “In summary, the next action I will perform is ```click [1234]```”.
5. Issue stop action when you think you have achieved the objective. Don’t generate anything after stop.